8. Express the roots of the equation z^3 − α^3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib.
(a) z^3 − 27 = 0
(b) z^3 + 8 = 0
(c) z^3 − i = 0.
1
Expert's answer
2020-05-26T18:51:15-0400

"z^3 −α^3 =0; \\
z^3-\alpha^3=(z-\alpha)(z^2+\alpha z+\alpha^2); \\
\implies (z-\alpha)(z^2+\alpha z+\alpha^2)=0; \\

 \implies z-\alpha=0 \ or \ z^2+\alpha z+\alpha^2=0; \\

\implies z_1=\alpha; D=\alpha^2-4\alpha^2=3\alpha^2;
\\

z_2=\dfrac{-\alpha+\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1+i\sqrt{3}}{2};

z_3=\dfrac{-\alpha-\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1-i\sqrt{3}}{2}" .

Thus roots of given equation are: "\alpha, \alpha w , \alpha w^2" .


a) Given equation is "z^3 - 27 =0" ;

By comparing with above equation we have "\alpha^3 = 27 \implies \alpha = 3"

So, roots of given equation are "3, 3 (\dfrac{-1+i\sqrt{3}}{2}), 3(\dfrac{-1-i\sqrt{3}}{2})" .


b) In this case "\alpha^3 = - 8 = (-2)^3 \implies \alpha = -2"

So, roots of given equation are "-2, -2 (\dfrac{-1+i\sqrt{3}}{2}), -2(\dfrac{-1-i\sqrt{3}}{2})" .


c) In this case, "\alpha^3 = i = (-i)^3 \implies \alpha = -i"

So, roots of given equation are "-i, -i (\dfrac{-1+i\sqrt{3}}{2}), -i(\dfrac{-1-i\sqrt{3}}{2})"

Thus, roots in form of a+ib are "-i, \dfrac{\sqrt{3}+i}{2}, \dfrac{-\sqrt{3}+i}{2}" .